OpenStudy (anonymous):

Match each item in Column A to an answer in Column B. Ignore the question mark. What property is shown by the equation? 1. (–22 + 17) + 13 = –22 + (17 + 13) 2. –(27 + (–3)) = –27 + 3 3. 35 + (–15 + 4) = (35 + (–15)) + 4 . –12 + 6 = 6 + (–12) Collum B A. commutative property B. associative property C. opposite of a sum property

OpenStudy (anonymous):

1. Associative property, because you move parentheses around 2. Opposite of a sum property, because the -'s turn to +'s and the +'s to -'s ...I can't read it anymore o.o all the equations got messed up

OpenStudy (anonymous):

3. Is also the associative property, again, moving parentheses

OpenStudy (anonymous):

4. Commutative property, because you're switching the order of -12 and 6

OpenStudy (anonymous):

Thanks can you answer some more please?

OpenStudy (anonymous):

Which is an example of the associative property? A. –(18 + 3) = –18 + (–3) B. –17 + 0 = –17 C. (–8 + 10) + 4. = –8 + (10 + 4) D. –5 + 7 = 7 + (–5)

OpenStudy (anonymous):

Why don't you do them and I'll check your answers?

OpenStudy (anonymous):

Yes please.

OpenStudy (anonymous):

The Answer is C, i belive.

OpenStudy (anonymous):

You're correct :)

OpenStudy (anonymous):

Alright, 3 more

OpenStudy (anonymous):

–(m + 15) = –18 + (–15)

OpenStudy (anonymous):

haha go for it

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

or 18

OpenStudy (anonymous):

Are you trying to figure out what m equals?

OpenStudy (anonymous):

Yeah. What value of m makes the statement true?

OpenStudy (anonymous):

ohh okay. Well, it doesn't look like they're applying the associative property because there are only 2 sums. And it's not the commutative property either, because the position of the 15 doesn't change. Which leaves which property?